### v1

# Radar-Guided Polynomial Fitting for Metric Depth Estimation

Abstract
We propose PolyRad, a novel radar-guided depth estimation method that introduces polynomial fitting to transform scaleless depth predictions from pretrained monocular depth estimation (MDE) models into metric depth maps. Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a simple yet fundamental insight: using polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust depth predictions non-uniformly across depth ranges. Although MDE models often infer reasonably accurate local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale-and-shift transformation insufficient given three or more of these regions. In contrast, PolyRad generalizes beyond linear transformations and is able to correct such misalignments by introducing inflection points. Importantly, our polynomial fitting framework preserves structural consistency through a novel training objective that enforces monotonicity via first-derivative regularization. PolyRad achieves state-of-the-art performance on the nuScenes, ZJU-4DRadarCam, and View-of-Delft datasets, outperforming existing methods by 30.3% in MAE and 37.2% in RMSE.
1 Introduction
Metric 3D reconstruction is critical for spatial tasks such as self-driving [34, 14], where it is necessary for one to perceive the structure of the 3D environment in order to navigate. In many such systems, multiple sensors—including cameras, lidar, and radar provide complementary information about the 3D scene. While lidar sensors offer dense, accurate point clouds, they are expensive and not widely available [39]. In contrast, radar sensors, particularly millimeter wave (mmWave) radars [20], return only about a hundred points per frame and are noisier [15]. Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles [10].
Figure 1: Once an MDE model predicts incorrect relative depths between three or more objects, a linear scale-and-shift (dashed) cannot resolve this misalignment. Instead, we learn a polynomial fit with inflection points, enabling flexible non-uniform corrections.
Unless camera baselines are known, images offer scaleless reconstruction. However, they are high-dimensional and sensitive to variations in illumination, object appearance, orientation, and camera viewpoint. Training robust models that can generalize across these factors often demands large-scale datasets, which are costly and time-intensive to collect. As a result, leveraging pretrained monocular depth estimation (MDE) models [41, 56, 4, 37] has emerged as a practical alternative. However, as these models only take in a single image, where reconstruction is inherently ill-posed, they typically infer scaleless relative depth, or metric depth that lacks the fidelity needed (see Tab. 2) for applications requiring accurate metric-scale reconstruction, e.g., navigation.
Existing approaches [60, 51, 58, 18, 8, 59] attempt to transform these MDE predictions into accurate metric depth maps using a global linear (scale-and-shift) transformation. This class of methods assumes that the reconstruction is off by a single scaling factor across the entire scene. While effective in estimating ordinal relationships within local regions, where MDE predictions typically exhibit reasonable local (object-level) reconstructions, such linear corrections fail when multiple objects are placed at incorrect depths relative to each other. Specifically, once an MDE model places three or more objects at incorrect relative depths, no simple scale-and-shift can reconcile this misalignment (see Fig. 1). Empirically, a learned linear fit on a challenging dataset such as nuScenes performs worse than even a deterministic estimator like median scaling (see Tab. 2), indicating susceptibility to outliers.
In this work, we challenge the common misconception that scale ambiguity in MDE is only up to an unknown global scale and shift. Instead, we propose to incorporate higher-order corrections through a polynomial transformation that allows for stretching and compressing at different depth levels, thereby enabling more flexible non-uniform adjustments. Our approach, PolyRad, transforms the predictions of pretrained MDE models by: learning prototypical patterns in the configuration of radar points, establishing spatial correspondences between radar and MDE features to encode a unified multi-sensor scene representation, and predicting polynomial coefficients that adaptively fit scaleless depth into metric depth. This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and inter-region misalignments. While one may adjust every prediction (pixel) of an MDE model [25], the number of degrees of freedom is the total number of pixels. Given radar point clouds of a few hundred points (many orders less than the number of pixels), this becomes an ill-posed problem. On the other hand, the number of degrees of freedom in polynomial expansion is limited to the number of polynomial terms, which we empirically found to be close to the cardinality of radar point clouds, allowing our solution to be better posed and more regular. To our knowledge, no prior work has explored polynomial fitting for adapting predictions of pretrained foundation models, making our approach a novel solution not only for radar-guided depth estimation but also for any task that leverages foundation models and refines their predictions using additional inputs as guidance.
However, learning such a flexible transformation is challenging because higher-order polynomials introduce a vast function space with many free parameters. Unlike a simple linear fit that relies on only scale and bias terms, polynomial fitting allows for numerous inflection points and nonlinearities. This expressiveness can inadvertently lead to harmful non-monotonic transformations—where the relative depth orderings of points are incorrectly reversed—if not properly constrained. We address this by introducing a regularization term that enforces the first derivative of the predicted metric depth with respect to the input MDE depth to generally remain close to 1, ensuring that incremental changes in the input MDE depth result in approximately equal changes in the predicted depth within local regions where MDE reconstructs structure with high fidelity. In essence, this regularizer encourages a smooth, monotonic transformation, mitigating unstable oscillations that can arise from overfitting high-degree polynomials while still enabling necessary corrections for cross-region misalignments.
Our contributions: We propose (1) a novel radar-guided depth estimation method that leverages polynomial fitting to transform scaleless MDE predictions into accurate metric depth maps. Our approach (2) introduces a fundamental insight: using polynomial coefficients predicted from radar points to enable non-uniform, depth-dependent corrections. We present (3) a principled geometric approach, where polynomial transformations naturally introduce inflection points that can correct misalignments between local areas that a global scale-and-shift transformation cannot. We design (4) a novel training objective that encourages monotonicity through first-derivative as regularization, ensuring structural consistency while allowing inter-region depth adjustments. Finally, (5) extensive experiments on the nuScenes, ZJU-4DRadarCam, and View-of-Delft datasets demonstrate that PolyRad achieves the state of the art, outperforming existing methods by 11.3%, 54.6%, and 35.2%, respectively.
2 Related Work
2.1 Monocular Depth Estimation
With recent advances in the scalability of neural networks [9, 5], monocular depth estimation (MDE) models can infer relative depth from a single image across diverse, unseen domains [22, 40, 55, 56, 37, 4], benefiting from large-scale datasets. Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation [16, 2], uncertainty estimation [38], optical flow [62], and visual odometry [46], to improve relative depth estimation. However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating absolute depth from a single image is an ill-posed problem. Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity [51]. One way to address this limitation is to incorporate range-sensing modalities such as lidar [21, 11], radar [45], or visual-inertial odometry [52].
2.2 Range-Camera Depth Estimation
Metric-scale depth estimation often requires recovering a dense depth map from sparse range measurements, guided by complementary camera inputs. Image-guided depth completion, such as lidar-camera depth estimation, leverages the strengths of each modality: images offer dense visual context and structural priors, while sparse depth provides absolute metric scale to resolve depth ambiguity. Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers [32, 33], late fusion, which maintains independent branches for different sensors before merging outputs [57, 54, 50, 12, 43], and multi-scale fusion, which captures both local details and global scene structure [23]. U-Net-like architectures with CSPN-based refinement [6] have been widely used for coarse-to-fine depth completion [17, 27, 36], with further improvements from adaptive convolutions [7], deformable convolutions [36, 53], and transformer-based attention mechanisms [42, 61].
2.3 Radar-Camera Depth Estimation
While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions [39]. In contrast, radar provides a cost-effective alternative [19], offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle [35, 47, 44]. The ubiquity of radar sensors [10] in existing automotive and robotic platforms further supports their integration into depth estimation pipelines. Leveraging radar for metric depth estimation not only reduces costs but also enhances the robustness and scalability of perception systems, making it an attractive choice for real-world deployment.
Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity [45] of radar point clouds. [26] uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map. [29] leverages Doppler velocity and optical flow to associate radar points with image pixels. Other methods refine radar depth using neighborhood pixel associations [30], height-extension of radar points to address elevation ambiguity [28], or by extracting self-supervisory signals from the radar point cloud [13]. [45] introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion. [24] mitigates distribution artifacts using sparse supervision, while [48] employs a two-stage confidence-driven approach. [49] uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features. Like our method, [25] starts with an initial MDE prediction but refines it through quasi-dense scale estimation, followed by local scale adjustment, rather than through a learned scene-fitting approach.
While these existing methods employ complex architectures to refine radar-derived spatial information, they all operate within the paradigm of completion or confidence-driven fusion. They often rely on multi-stage training, explicit radar-pixel association learning, or additional supervisory signals, greatly increasing model complexity. In contrast, PolyRad employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale. It bypasses the need for complex multi-stage systems that learn explicit correspondences while enabling flexible scene-adaptive depth corrections.
Figure 2: Overview of PolyRad. PolyRad transforms scaleless MDE predictions into metric depth using polynomial fitting guided by radar features. Learnable prototypes extract patterns in the configurations of radar point clouds and are used to compute aggregated radar features. The geometry-aware MDE features are fused with the radar features to yield a unified scene representation, which is used to predict coefficients for polynomial fitting, enabling non-uniform corrections that improve accuracy beyond simple scale-and-shift transformations.
3 Method Formulation
We aim to reconstruct a 3D scene from an RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a synchronized radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$ by fitting a scaleless depth map from a pretrained MDE foundation model to a metric depth map. Our method leverages MDE to seed scaleless depth predictions and estimates metric depth through polynomial fitting guided by cross-modal features that encode spatial correspondences between the 3D radar points and the dense scaleless depth predictions.
As MDE models are trained on millions of images, they are robust to typical visual nuisance variability, spanning from illumination changes and object appearances to viewpoint shifts. To circumvent the need to collect many large-scale synchronized radar-camera datasets, we begin with a frozen pretrained MDE model $M$, which serves as a learned geometric prior, to infer a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the image. We then encode the point cloud $C$ and the scaleless depth map $z$ separately. Through a fusion process that captures correspondences between them, we predict $N+1$ polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\cdots,\hat{c}_{N}\}$ that are used to transform $z$ into a metric depth map $\hat{d}$:
$\hat{d}\,=\,\hat{c}_{0}+\sum\nolimits_{i=1}^{N}\hat{c}_{i}z^{i}.$ (1)
3.1 Motivation
Existing methods [60, 51, 58, 18, 8, 59] refine MDE predictions by applying simple scale-and-shift transformations. However, this approach lacks the complexity needed to correct non-uniform variations from ground truth across different depths. A practical illustration of this shortcoming arises when multiple objects or local regions appear in the same scene: MDE may accurately infer the reconstruction within each object region, yet place them at incorrect relative depths with respect to each other. Such cross-object misalignments cannot be corrected by a single global scaling factor (see Fig. 1). This violates the assumption that the MDE reconstruction is up to an unknown global scale and shift.
To address this, our polynomial fitting approach exponentiates the MDE prediction to higher powers and learns coefficients that transform it into a metric depth map. Lower-order coefficients (including scale and shift) capture the global scene layout, while higher-order coefficients focus on local depth adjustments, correcting cross-object misalignments and fine-grained MDE errors.
An intuitive understanding of polynomial fitting can be drawn from the geometric perspective of depth transformations. A simple scale and shift operation, mathematically represented as $\hat{a}z+\hat{b}$, uniformly stretches or compresses the entire reconstruction by the same factor across all depth levels. While effective for coarse global corrections, such a linear transformation has zero inflection points, limiting its flexibility to address cross-region misalignments in relative depth predictions. In contrast, the total degree of our polynomial fitting method determines the maximum number of potential inflection points, where the curvature of the depth transformation can change. For a polynomial $f(z)=\sum_{i=0}^{N}\hat{c}_{i}\,z^{i}$, an inflection point $z^{*}$ occurs if
[…]
indicating where the second derivative changes sign. Our method learns inflection points to model transition regions where the curvature of the MDE error changes, allowing necessary non-uniform corrections.
nuScenes ZJU View-of-Delft
       Distance Method MAE RMSE MAE RMSE MAE RMSE
       50m Lin [26] [IROS ’20] 1598.2 3790.1 1576.3 3464.7 2894.9 5507.7
RadarNet [45] [CVPR ’23] 1727.7 3746.8 1430.5 3250.8 2944.8 6113.2
SparseBeatsDense [24] [ECCV ’24] 1524.5 3567.3 1424.4 3267.5 2909.9 5746.4
GET-UP [49] [WACV ’25] 1241.0 2857.0 1483.9 3220.5 2330.0 4565.0
RadarCam-Depth [25] [ICRA ’24] 1286.1 2964.3 1067.5 2817.4 1895.1 4458.7
PolyRad (Ours) 1146.7 2695.6 578.0 1108.6 1293.7 2420.6
       70m Lin [26] [IROS ’20] 1897.8 4558.7 1701.7 3853.3 3312.3 6480.5
RadarNet [45] [CVPR ’23] 2073.2 4590.7 1543.8 3655.3 3428.9 7331.9
SparseBeatsDense [24] [ECCV ’24] 1822.9 4303.6 1520.0 3593.4 3408.3 6914.7
GET-UP [49] [WACV ’25] 1541.0 3657.0 1651.5 3711.7 2758.0 5678.0
RadarCam-Depth [25] [ICRA ’24] 1587.9 3662.5 1157.0 3117.7 2095.4 4944.6
PolyRad (Ours) 1402.9 3269.7 603.7 1154.9 1442.9 2803.8
       80m Lin [26] [IROS ’20] 1988.4 4841.1 1739.5 3993.9 3468.1 6910.3
RadarNet [45] [CVPR ’23] 2179.3 4898.7 1578.4 3804.2 3597.0 7809.2
SparseBeatsDense [24] [ECCV ’24] 1927.0 4609.6 1548.4 3708.1 3584.0 7375.2
GET-UP [49] [WACV ’25] 1632.0 3932.0 1699.7 3882.6 2917.3 6145.1
RadarCam-Depth [25] [ICRA ’24] 1689.7 3948.0 1183.5 3229.0 2227.4 5385.8
PolyRad (Ours) 1454.1 3488.9 629.6 1171.3 1500.1 3951.8
Table 1: Quantitative results on nuScenes, ZJU-4DRadarCam, and View-of-Delft test sets evaluated with various maximum depths.
3.2 Radar Processing
We begin with a radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$, where $N_{C}$ denotes the number of points and each point is represented using its coordinates in three-dimensional camera space. Radar point clouds, while noisy and sparse, provide metric depth measurements. We encode $C$ using a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
To facilitate effective radar feature aggregation, we introduce a set of learnable prototypes $P\in\mathbb{R}^{N_{P}\times c_{r}}$, where these prototypes can be interpreted as capturing diverse spatial and geometric properties present in the radar point cloud. Each prototype learns to focus on specific patterns in the configuration of radar points, making them highly expressive and adaptable across varying scenes. Unlike existing works [45, 25] that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud. This allows for selective feature aggregation that mitigates the impact of outliers, resulting in more robust depth predictions.
The next step is to match the learnable prototypes to the most relevant radar features for a given input. We allow each prototype to dynamically attend to the radar points that provide the most relevant spatial and geometric signals, thereby aggregating a sparse point cloud into a global scene-descriptive representation. To achieve this, we project the point-wise radar features $F_{r}$ into a key-value space, where they can be efficiently retrieved by the prototypes $P$. We compute the aggregated radar features $R$:
$R=\text{softmax}\left(\frac{P\times(\psi_{r}^{K}(F_{r}))^{T}}{\sqrt{c_{r}}}\right)\psi_{r}^{V}(F_{r}),$ (3)
where $\psi_{r}^{K}$ and $\psi_{r}^{V}$ are MLPs that project the point-wise radar features $F_{r}$. Comparatively, applying depth completion methods that densify a sparse projection of points using surrounding context [31, 36] to a radar point cloud results in poor performance (see Supp. Mat). This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution [45].
Next, the scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the MDE model is encoded using a depth encoder, $f_{z}$, to produce depth features $Z\in\mathbb{R}^{(H\times W)\times c_{z}}$. We capitalize on the invariants learned through large-scale MDE training—such as robustness to color variations, illumination changes, and diverse object poses—by extracting features from the MDE depth maps, which leads to features encoding shapes and surfaces. We note that variations in the shape of objects or their geometry are comparatively smaller than variations in visual appearance, which are highly sensitive to transient factors such as perspective and lighting. We then learn spatial correspondences between the depth features $Z$ and the radar features $R$ to produce a unified scene representation that fuses the structural information encoded in the MDE predictions with the metric cues provided by the radar features. An important insight is that our depth encoder $f_{z}$ yields more stable geometric context by matching radar point configurations to shapes and surfaces that remain relatively consistent across scenes, rather than to pixel-wise color intensities that can vary arbitrarily with lighting and viewpoint.
To this end, we fuse the depth features $Z$ and the aggregated radar features $R$ to yield a unified scene representation
$S=\text{softmax}\left(\frac{Z\times(\psi_{R}^{K}(R))^{T}}{\sqrt{c_{r}}}\right)\psi_{R}^{V}(R),$ (4)
where $\psi_{R}^{K}$ and $\psi_{R}^{V}$ are MLPs that project the aggregated radar features $R$. The fused representation $S$ is passed through a shallow convolutional neural network (CNN) $f_{s}$ followed by a global average pooling (GAP) layer to yield
$\bar{S}=\text{GAP}(f_{s}(S)),$ (5)
which is a $c_{s}$-dimensional feature vector. The GAP layer aggregates spatial information from the entire scene, ensuring that $\bar{S}$ captures global context. This final scene representation $\bar{S}$ is then fed into an MLP $\psi_{s}$, which predicts the $N+1$ polynomial coefficients as a vector $\hat{c}$:
$\hat{c}=\psi_{s}(\bar{S})\in\mathbb{R}^{N+1}.$ (6)
These coefficients allow the model to adaptively refine the initial depth predictions from the MDE model, with each coefficient adjusting depth at different scales and granularities. Lower-order terms capture broad scene structure, while higher-order terms enable fine-grained and inter-region corrections, resulting in a high-fidelity metric-scale depth map.
Figure 3: Qualitative comparison on nuScenes. Baseline methods GET-UP and RadarCam-Depth (RC-D) fail to correctly reconstruct the 3D scene, resulting in objects with incorrect or missing depth predictions. While raw MDE outputs reasonable scaleless reconstructions, several cross-object misalignments are apparent. In contrast, PolyRad leverages polynomial fitting to infer high-fidelity depth predictions.
3.3 Polynomial Fitting
Given a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ predicted by the MDE foundation model, our goal is to transform it into a metric depth map $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ using the polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\dots,\hat{c}_{N}\}$ predicted by our network (Sec. 3.2).
Formally, we express the final depth map $\hat{d}$ as:
$\begin{split}\hat{d}(x,y)=\sum_{i=0}^{N}\,\hat{c}_{i}\cdot z(x,y)^{i},\;\;\forall(x,y)\in H\times W\end{split}$ (7)
This polynomial formulation applies the learned coefficients to progressively higher powers of the scaleless depth map, enabling complex, non-linear transformations that go beyond simple scale-and-shift adjustments. Each pixel’s final depth is thus computed by combining multiple weighted terms, where each term corresponds to a different polynomial order of the initial depth prediction.
This is in contrast to directly mapping to a metric-scale depth map [45, 25], where the degrees of freedom comprise every predicted pixel. Each corresponding point and pixel prediction can be viewed as a constraint or equation in our estimation system. As the radar point cloud (hundreds of points) only occupies a small subset of the image space, the solution is underdetermined and often irregular (see Fig. 3). Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab. 3). PolyRad offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
In terms of overhead, compared to predicting scale and shift coefficients for a linear transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost: an MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Supp. Mat).
Geometric Intuition. Applying polynomial fitting to the scaleless depth map can be viewed as a flexible, depth-dependent correction mechanism. In contrast to a single scale-and-shift operation, which uniformly adjusts the entire depth field by a single global factor, polynomial transformations introduce multiple degrees of freedom—higher-order terms enable correction of misalignments in relative depth between objects and local regions.
By introducing polynomial terms, our method enables depth adjustments that vary according to the initial depth estimates. Lower-order terms ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) capture broad scene structure, ensuring that near and far objects maintain reasonable depth relationships. Higher-order terms ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) refine depth predictions by introducing localized adjustments that correct subtle variations, such as mid-range depth discrepancies or fine object boundaries.
The concept of inflection points provides an intuitive understanding of this approach. Each additional polynomial order introduces more potential inflection points, allowing the depth transformation to shift curvature where needed. This flexibility enables “stretching” or “compressing” different depth levels: areas already near their correct metric depth receive minimal adjustment, while regions suffering larger errors (e.g., due to misalignment with other regions) undergo more substantial correction. Higher-order terms magnify small discrepancies in the initial MDE predictions, making them more apparent to the model and easier to correct. In this way, the degree of the polynomial, selected as a hyperparameter, directly reflects the capacity of our model to perform non-uniform depth-dependent adjustments.
The sign of the predicted polynomial coefficients furthers our interpretation. Positive coefficients for higher-order terms push depth values outward while negative ones pull them inward, effectively expanding or contracting selected depth intervals to correct local errors. In doing so, these coefficients shape the curvature of the polynomial, dictating where and how inflection points arise. For instance, if the model consistently overestimates distant objects, a negative high-order coefficient can compress that region, whereas a positive coefficient might be learned to correct underestimations. By treating coefficient signs as dynamic anchors for curvature changes, the polynomial fitting framework provides an interpretable mechanism (see Fig. 4) for how our method adjusts depth predictions at different scales.
Figure 4: PolyRad leverages higher-degree polynomials to reduce error by correcting depth misalignments using inflection points.
3.4 Loss Function
We employ a loss function comprising three terms to guide the learning of the polynomial coefficients and ensure accurate depth estimation. Our loss is defined as:
$\mathcal{L}=\lambda_{1}\|\hat{d}-d_{\text{gt}}\|_{1}+\lambda_{2}\|\hat{d}-d_{\text{gt}}\|_{2}^{2}+\lambda_{3}\left\|1-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (8)
where $d_{\text{gt}}\in\mathbb{R}_{+}^{H\times W}$ is the ground truth metric depth map, $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ is our predicted metric depth, and $z\in\mathbb{R}_{+}^{H\times W}$ is the scaleless depth map predicted by the MDE model.
The first two terms, $L_{1}$ and $L_{2}$ losses, ensure that our predicted depth map $\hat{d}$ closely matches the ground truth $d_{\text{gt}}$ by penalizing discrepancies. The $L_{1}$ term is less sensitive to outliers and thus promotes robustness, while the $L_{2}$ term penalizes larger errors more significantly.
The novel component of our loss function lies in the third term, which encourages the first derivative of the predicted depth map $\hat{d}$ with respect to the input scaleless depth $z$ to generally be close to 1. Formally, we define:
$\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}(x,y)=\sum_{i=1}^{N}\,i\,\hat{c}_{i}\,z(x,y)^{i-1}.$ (9)
This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression [1]. Generally, within an object or local region, a pixel with a higher initial scaleless depth value should not be assigned a lower final metric depth value compared to a pixel with a lower initial scaleless depth. The preservation of ordinal relationships is an inductive bias grounded in the assumption that the MDE model provides a reasonably accurate estimation of intra-object relative depth.
However, polynomial fitting introduces a large function space, with expressiveness increasing as the polynomial degree $N$ grows, potentially disrupting local monotonic depth ordering if not properly constrained. Our regularization term addresses this by encouraging that, on average across the scene, the slope of our polynomial transformation remains close to 1, thus preserving monotonicity within local regions. At the same time, since there are fewer measurements between objects, PolyRad is free to correct cross-object misalignments while preserving local depth ordinality.
A first derivative close to 1 means that for every incremental change in the input scaleless depth $z$, there is an approximately equal incremental change in the predicted metric depth $\hat{d}$. This constraint ensures that our model does not overfit to noisy radar data or introduce spurious corrections by avoiding highly non-linear transformations that could disrupt the monotonicity of depth within local regions. Furthermore, this regularization term mitigates oscillatory polynomial curves, a known challenge when using higher-degree polynomials [3]. Without this constraint of regularity, the polynomial may excessively oscillate due to overfitting, leading to erratic depth predictions in some regions.
4 Experiments
Baseline Methods. We consider five recent radar-camera depth estimation baselines: RadarNet [45], Lin [26], SparseBeatsDense [24], GET-UP [49], and RadarCam-Depth [25]. Notably, RadarCam-Depth also leverages radar points to rescale initial scaleless MDE predictions.
Datasets. We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Supp. Mat for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
4.1 Main Results
PolyRad reduces MAE by 12.1% and RMSE by 10.5% on nuScenes, 46.8% and 62.4% on ZJU, and 31.8% and 38.5% on VoD, achieving SOTA results for all datasets (see Tab. 1).
We qualitatively compare PolyRad against all baseline methods on the challenging nuScenes dataset, providing a visual demonstration of its improved depth estimation performance (see Fig. 3). PolyRad more accurately predicts metric scale and exhibits fewer global misalignments. Furthermore, it is able to extrapolate depth predictions of object surfaces from learned photometric priors, an ability inherited from the backbone MDE model (UniDepth) trained on large-scale datasets. Our polynomial fitting approach then non-uniformly corrects depth discrepancies across the scene, leading to significant performance improvement over initial MDE predictions. Introducing multiple degrees of freedom enables PolyRad to correct misalignments of objects relative to each other, improving accuracy by correcting consistent over- or under-estimations in local regions. For instance, the raw MDE prediction for Image C incorrectly places the bus stop roof at around the same depth as the more distant tree branch, as highlighted on the right. Additionally, MDE does not accurately infer the boundary between the curb and asphalt, again placing both at erroneously similar depths. With a learned polynomial transformation, PolyRad corrects these misalignments with depth-dependent adjustments, accurately separating the curb from the asphalt and placing the bus stop roof and tree branch at their correct depths.
While RadarCam-Depth also utilizes an MDE backbone, its complex processing of raw MDE predictions can be seen to distort initially correct depth structure. In Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths. Furthermore, overfitting to specific regions within a scene—particularly those with denser ground truth—may reduce overall error relative to the MDE model but can lead to unintended geometric artifacts. This can result in the omission of entire structures in the predicted depth map, such as the highlighted building on the left of Image A and the construction crane arms in Image B. GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects across all three samples: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
nuScenes
Method MAE RMSE
DPT [41] 5188.2 6884.5
Depth Anything [56] 2404.9 4851.1
Depth Pro [4] 3835.0 6600.3
UniDepth [37] 2129.8 4887.7
RadarCam-Depth w/ DPT 1689.7 3948.0
RadarCam-Depth w/ Depth Anything 1953.6 5107.8
RadarCam-Depth w/ Depth Pro 3417.1 6462.0
RadarCam-Depth w/ UniDepth 1872.0 4321.2
PolyRad w/ DPT 1525.6 3745.0
PolyRad w/ Depth Anything 1515.1 3719.4
PolyRad w/ Depth Pro 1627.7 4143.9
PolyRad w/ UniDepth 1454.1 3488.9
Table 2: MDE backbone comparative studies on nuScenes.
4.2 Comparative Studies
MDE Models. We evaluate PolyRad using four monocular depth estimation (MDE) backbones: DPT [41], Depth Anything [56], Depth Pro [4], and UniDepth [37]. DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see Tab. 2 and Supp. Mat), and thus we process them the same way as scaleless depth.
Tab. 2 compares raw MDE performance to their performance when used as backbones for RadarCam-Depth and PolyRad. Raw MDE predictions show large errors due to scale ambiguity, while RadarCam-Depth provides moderate improvements. In contrast, PolyRad consistently reduces error across all backbones, demonstrating the effectiveness of polynomial fitting. For DPT and Depth Anything, outputs are first inverted when used as a backbone, and are further median-scaled for the reported raw results. Among all configurations, PolyRad w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 30.2% and over RadarCam-Depth w/ UniDepth by 20.8%.
Polynomial Degree. As polynomial degree—selected as a hyperparameter—increases, so does the number of inflection points, allowing PolyRad to correct misalignments between the many scene elements in Fig. 4, including cars, trucks, trees, buildings, and traffic lights. This is reflected in the successively improved error maps. Tab. 3 quantifies the performance gains achieved with higher-degree polynomials. At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
Ablations. We evaluate the impact of key components of PolyRad in Tab. 4. Replacing learnable prototypes with self-attention on the radar point features degrades performance, demonstrating that prototypes capture relevant patterns within radar point configurations. Removing cross-modality fusion results in the largest performance drop, highlighting the necessity of leveraging correspondences between MDE and radar features. Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
nuScenes ZJU
Polynomial Degree MAE RMSE MAE RMSE
1 (Scale + Shift) 2156.8 4491.3 1078.2 2405.5
2 1715.2 3840.6 901.0 1822.7
4 1482.7 3510.1 791.2 1516.4
6 1466.9 3496.0 670.3 1297.9
8 1454.1 3488.9 629.6 1171.3
10 1463.7 3494.5 643.3 1184.5
Table 3: Sensitivity study of polynomial degree hyperparameter.
nuScenes VoD
Ablated Component MAE RMSE MAE RMSE
learnable prototypes 1615.5 3629.0 1619.3 4162.7
cross-modality att. 2238.8 4817.4 2147.9 4926.1
monotonicity loss 1921.1 4399.5 1924.5 4660.3
no ablations 1454.1 3488.9 1500.1 3951.8
Table 4: Ablation studies of architecture and loss components.
5 Discussion
Limitations. PolyRad requires tuning the polynomial degree as a hyperparameter, which may vary across datasets. Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
Summary. We are the first to formulate radar-camera depth estimation as a scene-fitting problem, leveraging high-degree polynomials to transform MDE predictions into accurate metric depth. Our principled and efficient approach demonstrates that a fundamental insight—shifting from simple scale-and-shift to flexible, radar-guided polynomial transformations—can outperform more complex systems.